# Profinite rigidity witnessed by Dehn fillings of cusped hyperbolic 3-manifolds

### 03 Mar 2026 v3

@@ 6. Profinite rigidity from exceptional slopes
6.1. Some Dehn-surgeries on the Whitehead link
@@ 6.3. More hyperbolic knot complements in $S^{3}$
    $\mathcal{E}_{t}(\mathcal{K}_{-1})=\{-4,0,4\}$. In particular, $\Delta(-4,4)=8$
@@
    For $m,n\in\mathbb{Z}\setminus\{-1,0,1\}$, $\widehat{\pi_{1}X_{\mathcal{J}_{m}}}\cong\widehat{\pi_{1}X_{\mathcal{J}_{n}}}$ if and only if $m=\pm n$; in this case, $\mathcal{J}_{m}$ and $\mathcal{J}_{n}$ are either equivalent or the mirror images of each other.
@@
Similarly, if $\widehat{\pi_{1}X_{\mathcal{J}_{m}}}\cong\widehat{\pi_{1}X_{\mathcal{J}_{n}}}$, then $\{-2-9m,2-9m\}=\mathcal{E}_{t}(\mathcal{J}_{m})=\pm\mathcal{E}_{t}(\mathcal{J}_{n})=\pm\{-2-9n,2-9n\}$ by Proposition 6.10, which implies $m=\pm n$. In addition, $\mathcal{J}_{n}$ can be viewed as a $\frac{1}{n}$-surgery on the link $\mathcal{L}$, while $\mathcal{L}$ is achiral, i.e. equivalent to its mirror image. Thus, the $\frac{1}{n}$-surgery on $\mathcal{L}$ yields the mirror-image of the $\frac{1}{n}-$surgery on $\mathcal{L}$. This implies that $\mathcal{J}_{n}$ and $\mathcal{J}_{-n}$ are the mirror images of each other. In particular, $X_{\mathcal{J}_{n}}$ is homeomorphic to $X_{\mathcal{J}_{-n}}$, so $\widehat{\pi_{1}X_{\mathcal{J}_{n}}}\cong\widehat{\pi_{1}X_{\mathcal{J}_{-n}}}$. This finishes the proof of (2).
@@
For brevity of notation, let $K$ denote one of the above listed knots, i.e. $K$ belongs to $\mathcal{K}_{n}$ ($n\in\mathbb{Z}\setminus\{0,1\}$), $\mathcal{J}_{n}$} ($n\in\mathbb{Z}\setminus\{-1,0,1\}$), and K(3,1,n,0){K(3,1,-n,0)} ($n\in\mathbb{Z}\setminus\{0\}$), and we denote $M=X_{K}$.
@@
When $K=K(3,1,n,0)$, Proposition 6.10 (4) implies that there exist $c_{1},c_{2}\in\mathcal{E}_{t}(K)$ such that $\Delta(c_{1},c_{2})=5$. The above reasoning shows that there exist $c_{1}^{\prime},c_{2}^{\prime}\in\mathcal{E}_{t}(K^{\prime})$ such that $\Delta(c_{1}^{\prime},c_{2}^{\prime})=5$. According to Proposition 6.9 (3), $K^{\prime}$ is either equivalent or mirror image to a Eudave-Muñoz knot $K(3,1,n^{\prime},0)$ ($n^{\prime}\in\mathbb{Z}\setminus\{0\}$). Thus, $\widehat{\pi_{1}M}\cong\widehat{\pi_{1}N}$ implies $\widehat{\pi_{1}X_{K(3,1,n,0)}}\cong\widehat{\pi_{1}X_{K(3,1,n^{\prime},0)}}$. Lemma 6.11 then guarantees $n=n^{\prime}$, i.e. $N\cong X_{K^{\prime}}\cong X_{K(3,1,n,0)}\cong M$.
@@ 6.5. Detecting Klein bottles and Dehn surgeries on $\mathcal{L}$ and $L_{B}$
Recall that $\mathcal{L}$ is the link shown in Figure 1(c). Similarly, let $\mathcal{L}(r)$ denote $r$-surgery on one component of $\mathcal{L}$, and $\mathcal{L}(r,s)$ denote $r$-surgery on one component and $s$-surgery on the other component of $\mathcal{L}$. Note that $\mathcal{L}$ is achiral, i.e. equivalent to its mirror image. Thus,
@@
Recall the notations that $\mathcal{N}$ is the orientable $I$-bundle over Klein bottle, and $M_{1}$ is the complement of the Whitehead link $\mathcal{W}$ in $S^{3}$. In the following context, we denote $M_{3_{1}}$ as the complement of the trefoil in $S^{3}$, i.e. $M_{3_{1}}$ is a Seifert fibered space $(D^{2};(2,1),(3,1))$.
@@
    $L_{B}(0,0)\cong L_{B}(0,4)$ can be decomposed as $\mathcal{N}\cup_{T^{2}}M_{1}$, i.e. gluing $\mathcal{N}$ and $M_{1}$ along a torus boundary.
@@
(2) According to (6.15), $L_{B}(0,r,0)\cong L_{B}(0,0,r)$ and $L_{B}(0,r,4)\cong L_{B}(0,4,r)$. Thus, they can be viewed as Dehn fillings on $L_{B}(0,0)$ and $L_{B}(0,4)$. According to (1), $L_{B}(0,0,r)$ and $L_{B}(0,4,r)$ can be decomposed by an embedded torus into $\mathcal{N}\cup M^{\prime}$, where $M^{\prime}$ is a Dehn filling of $M_{1}$ on one of its boundary components $\partial_{1}M_{1}$. The Dehn fillings of $M_{1}$ on $\partial_{1}M_{1}$ have been classified by [37, Table A.1]. In fact, there is exactly one slope yielding the solid torus $D^{2}\times S^{1}$, while the other slopes yield either a hyperbolic manifold, a Seifert fibered space with hyperbolic base orbifold, or a graph manifold with non-trivial JSJ-decomposition. In other words, there is only one element $r_{0}\in\mathbb{Q}\cup\{\infty\}$ such that $L_{B}(0,0,r_{0})\cong\mathcal{N}\cup_{T^{2}}(D^{2}\times S^{1})$, which is indeed non-aspherical; while for any $r\neq r_{0}\in\mathbb{Q}\cup\{\infty\}$, $L_{B}(0,0,r)\cong\mathcal{N}\cup_{T^{2}}M^{\prime}$ is aspherical. We claim that $r_{0}=\infty$. In fact, by (6.15), $L_{B}(0,0,\infty)\cong L_{B}(0,\infty,0)$ is a Dehn filling on $L_{B}(0,\infty)$. [32, Lemma 2.3 (1)] showed that $L_{B}(0,\infty)\cong\mathcal{N}$, and thus $L_{B}(0,\infty,0)\cong\mathcal{N}\cup_{T^{2}}(D^{2}\times S^{1})$. Similarly, there is only one element $r_{4}\in\mathbb{Q}\cup\{\infty\}$ such that $L_{B}(0,4,r_{4})\cong\mathcal{N}\cup_{T^{2}}(D^{2}\times S^{1})$ is spherical, while for any $r\neq r_{4}\in\mathbb{Q}\cup\{\infty\}$, $L_{B}(0,4,r_{4})\cong\mathcal{N}\cup_{T^{2}}M^{\prime}$ is aspherical. One can similarly show that $r_{4}=\infty$ by [32, Lemma 2.3 (1)]. Thus, this proves (2).
@@
We have shown by Corollary 6.19 that $\max\{\Delta(c_{1},c_{2})\mid c_{1},c_{2}\in\mathcal{E}_{k}(M,\partial M)\}\geq 4$. If $\max\{\Delta(c_{1},c_{2})\mid c_{1},c_{2}\in\mathcal{E}_{k}(M,\partial M)\}>4$, then [33, Corollary 1.5] shows that $M$ is homeomorphic to one of $\mathcal{W}(5),\mathcal{W}(-1),\mathcal{W}(-2),\mathcal{W}(-4)$. Thus, it suffices to show that $M$ is not homeomorphic to $\mathcal{W}(5),\mathcal{W}(-1),\mathcal{W}(-2),\mathcal{W}(-4)$.
@@
(3) For each hyperbolic vertex $v\in V(\Gamma_{M})$, $f_{v}:\widehat{\pi_{1}M_{v}}\stackrel{{\scriptstyle\cong}}{{\longrightarrow}}\widehat{\pi_{1}N_{f^{\dagger}(v)}}$ respects the peripheral structure according to [56, Lemma 6.3]. In fact, according to Proposition A.3 (4), $f_{v}$ matches up the boundary components of $M_{v}$ coming from the JSJ-tori (i.e. corresponding to the edges $e\in E(\Gamma_{M})$ adjoining $v$) to the boundary components of $N_{f^{\dagger}(v)}$ coming from the JSJ-tori (i.e. corresponding to the edges $f^{\dagger}(e)$). Thus, $f_{v}$ matches up the remaining boundary components, i.e. matches up $M_{v}\cap\partial M$ with $N_{f^{\dagger}(v)}\cap\partial N$. Therefore, up to reordering, $f_{v}$ sends $\overline{\pi_{1}\partial_{i}^{hyp}M}$ to a conjugate of $\overline{\pi_{1}\partial_{i}^{hyp}N}$ in $\widehat{\pi_{1}N_{f^{\dagger}(v)}}$. The commutative diagram in Proposition A.3 (3) then implies that $f$ in fact sends $\overline{\pi_{1}\partial_{i}^{hyp}M}$ to a conjugate of $\overline{\pi_{1}\partial_{i}^{hyp}N}$ in $\widehat{\pi_{1}N}$.

### 16 Mar 2026 v4

@@ 6. Profinite rigidity from exceptional slopes
6.1. Some Dehn surgeries on the Whitehead link
@@ 6.3. More hyperbolic knot complements in $S^{3}$
    $\mathcal{E}_{t}(\mathcal{K}_{-1})=\{-4,0,4\}$. In particular, $\Delta(-4,4)=8$.
@@
    For $m,n\in\mathbb{Z}\setminus\{-1,0,1\}$, $\widehat{\pi_{1}X_{\mathcal{J}_{m}}}\cong\widehat{\pi_{1}X_{\mathcal{J}_{n}}}$ if and only if $m=\pm n$; in this case, $\mathcal{J}_{m}$ and $\mathcal{J}_{n}$ are either isotopic or the mirror images of each other.
@@
Similarly, if $\widehat{\pi_{1}X_{\mathcal{J}_{m}}}\cong\widehat{\pi_{1}X_{\mathcal{J}_{n}}}$, then $\{-2-9m,2-9m\}=\mathcal{E}_{t}(\mathcal{J}_{m})=\pm\mathcal{E}_{t}(\mathcal{J}_{n})=\pm\{-2-9n,2-9n\}$ by Proposition 6.10, which implies $m=\pm n$. In addition, $\mathcal{J}_{n}$ can be viewed as a $\frac{1}{n}$-surgery on the link $\mathcal{L}$, while $\mathcal{L}$ is achiral, ie isotopic to its mirror image. Thus, the $\frac{1}{n}$-surgery on $\mathcal{L}$ yields the mirror-image of the $\frac{1}{n}-$surgery on $\mathcal{L}$. This implies that $\mathcal{J}_{n}$ and $\mathcal{J}_{-n}$ are the mirror images of each other. In particular, $X_{\mathcal{J}_{n}}$ is homeomorphic to $X_{\mathcal{J}_{-n}}$, so $\widehat{\pi_{1}X_{\mathcal{J}_{n}}}\cong\widehat{\pi_{1}X_{\mathcal{J}_{-n}}}$. This finishes the proof of (2).
@@
For brevity of notation, let $K$ denote one of the above listed knots, ie $K$ belongs to $\mathcal{K}_{n}$ ($n\in\mathbb{Z}\setminus\{0,1\}$), $\mathcal{J}_{n}$} ($n\in\mathbb{Z}\setminus\{-1,0,1\}$), and K(3,1,n,0){K(3,1,-n,0)} ($n\in\mathbb{Z}\setminus\{0\}$), and we denote $M=X_{K}$.
@@
When $K=K(3,1,n,0)$, Proposition 6.10 (4) implies that there exist $c_{1},c_{2}\in\mathcal{E}_{t}(K)$ such that $\Delta(c_{1},c_{2})=5$. The above reasoning shows that there exist $c_{1}^{\prime},c_{2}^{\prime}\in\mathcal{E}_{t}(K^{\prime})$ such that $\Delta(c_{1}^{\prime},c_{2}^{\prime})=5$. According to Proposition 6.9 (3), $K^{\prime}$ is either equivalent or mirror image to a Eudave-Muñoz knot $K(3,1,n^{\prime},0)$ ($n^{\prime}\in\mathbb{Z}\setminus\{0\}$). Thus, $\widehat{\pi_{1}M}\cong\widehat{\pi_{1}N}$ implies $\widehat{\pi_{1}X_{K(3,1,n,0)}}\cong\widehat{\pi_{1}X_{K(3,1,n^{\prime},0)}}$. Lemma 6.11 then guarantees $n=n^{\prime}$, ie $N\cong X_{K^{\prime}}\cong X_{K(3,1,n,0)}\cong M$.
@@ 6.5. Detecting Klein bottles and Dehn surgeries on $\mathcal{L}$ and $L_{B}$
Recall that $\mathcal{L}$ is the link shown in Figure 1(c). Similarly, let $\mathcal{L}(r)$ denote $r$-surgery on one component of $\mathcal{L}$, and $\mathcal{L}(r,s)$ denote $r$-surgery on one component and $s$-surgery on the other component of $\mathcal{L}$. Note that $\mathcal{L}$ is achiral, ie equivalent to its mirror image. Thus,
@@
Recall the notations that $\mathcal{N}$ is the orientable $I$-bundle over Klein bottle, and $M_{1}$ is the complement of the Whitehead link $\mathcal{W}$ in $S^{3}$. In the following context, we denote $M_{3_{1}}$ as the complement of the trefoil in $S^{3}$, ie $M_{3_{1}}$ is a Seifert fibered space $(D^{2};(2,1),(3,1))$.
@@
    $L_{B}(0,0)\cong L_{B}(0,4)$ can be decomposed as $\mathcal{N}\cup_{T^{2}}M_{1}$, ie gluing $\mathcal{N}$ and $M_{1}$ along a torus boundary.
@@
(2) According to (6.15), $L_{B}(0,r,0)\cong L_{B}(0,0,r)$ and $L_{B}(0,r,4)\cong L_{B}(0,4,r)$. Thus, they can be viewed as Dehn fillings on $L_{B}(0,0)$ and $L_{B}(0,4)$. According to (1), $L_{B}(0,0,r)$ and $L_{B}(0,4,r)$ can be decomposed by an embedded torus into $\mathcal{N}\cup M^{\prime}$, where $M^{\prime}$ is a Dehn filling of $M_{1}$ on one of its boundary components $\partial_{1}M_{1}$. The Dehn fillings of $M_{1}$ on $\partial_{1}M_{1}$ have been classified by [37, Table A.1]. In fact, there is exactly one slope yielding the solid torus $D^{2}\times S^{1}$, while the other slopes yield either a hyperbolic manifold, a Seifert fibered space with hyperbolic base orbifold, or a graph manifold with non-trivial JSJ-decomposition. In other words, there is only one element $r_{0}\in\mathbb{Q}\cup\{\infty\}$ such that $L_{B}(0,0,r_{0})\cong\mathcal{N}\cup_{T^{2}}(D^{2}\times S^{1})$, which is indeed non-aspherical; while for any $r\neq r_{0}\in\mathbb{Q}\cup\{\infty\}$, $L_{B}(0,0,r)\cong\mathcal{N}\cup_{T^{2}}M^{\prime}$ is aspherical. We claim that $r_{0}=\infty$. In fact, by (6.15), $L_{B}(0,0,\infty)\cong L_{B}(0,\infty,0)$ is a Dehn filling on $L_{B}(0,\infty)$. [32, Lemma 2.3 (1)] showed that $L_{B}(0,\infty)\cong\mathcal{N}$, and thus $L_{B}(0,\infty,0)\cong\mathcal{N}\cup_{T^{2}}(D^{2}\times S^{1})$ is non-aspherical. Similarly, there is only one element $r_{4}\in\mathbb{Q}\cup\{\infty\}$ such that $L_{B}(0,4,r_{4})\cong\mathcal{N}\cup_{T^{2}}(D^{2}\times S^{1})$ is non-aspherical, while for any $r\neq r_{4}\in\mathbb{Q}\cup\{\infty\}$, $L_{B}(0,4,r_{4})\cong\mathcal{N}\cup_{T^{2}}M^{\prime}$ is aspherical. One can similarly show that $r_{4}=\infty$ by [32, Lemma 2.3 (1)]. Thus, this proves (2).
@@
We have shown by Corollary 6.19 that $\max\{\Delta(c_{1},c_{2})\mid c_{1},c_{2}\in\mathcal{E}_{k}(M,\partial M)\}\geq 4$. If $\max\{\Delta(c_{1},c_{2})\mid c_{1},c_{2}\in\mathcal{E}_{k}(M,\partial M)\}>4$, then [33, Corollary 1.5] shows that $M$ is homeomorphic to one of $\mathcal{W}(5),\mathcal{W}(-1),\mathcal{W}(-2),\mathcal{W}(-4)$. We verify that $M$ is not homeomorphic to $\mathcal{W}(5),\mathcal{W}(-1),\mathcal{W}(-2),\mathcal{W}(-4)$.
@@
(3) For each hyperbolic vertex $v\in V(\Gamma_{M})$, $f_{v}:\widehat{\pi_{1}M_{v}}\stackrel{{\scriptstyle\cong}}{{\longrightarrow}}\widehat{\pi_{1}N_{f^{\dagger}(v)}}$ respects the peripheral structure according to [56, Lemma 6.3]. In fact, according to Proposition A.3 (4), $f_{v}$ matches up the boundary components of $M_{v}$ coming from the JSJ-tori (ie corresponding to the edges $e\in E(\Gamma_{M})$ adjoining $v$) to the boundary components of $N_{f^{\dagger}(v)}$ coming from the JSJ-tori (ie corresponding to the edges $f^{\dagger}(e)$). Thus, $f_{v}$ matches up the remaining boundary components, ie matches up $M_{v}\cap\partial M$ with $N_{f^{\dagger}(v)}\cap\partial N$. Therefore, up to reordering, $f_{v}$ sends $\overline{\pi_{1}\partial_{i}^{hyp}M}$ to a conjugate of $\overline{\pi_{1}\partial_{i}^{hyp}N}$ in $\widehat{\pi_{1}N_{f^{\dagger}(v)}}$. The commutative diagram in Proposition A.3 (3) then implies that $f$ in fact sends $\overline{\pi_{1}\partial_{i}^{hyp}M}$ to a conjugate of $\overline{\pi_{1}\partial_{i}^{hyp}N}$ in $\widehat{\pi_{1}N}$.